OpenStudy (xoxolove456):

Help with this world problem !!!! Cole and Mary can clean a room together in 5 hours. Working alone, Cole can clean the room in 1 hour less time than Mary can. How long does it take Mary to clean the room alone ?????!

OpenStudy (nonstopper):

5cole+5mary=x ycole =x ---- cole takes y hours to clean the room (y+1)mary=x ---- mary takes y+1 hours to clean the room Re-arrange the latter two: cole=x/y mary=x/(y+1) Substitute into the original: (5x/y)+(5x/y-1)=x divide by 5x-> 1/y+(1/(y-1))=1/5 solve for y, y =10.52 cole takes 10.52 hours mary takes 11.52 hours I'm not sure, but that's my best go. Good question.

OpenStudy (nonstopper):

mistake on line 8, it's 5x/(y+1). the answer is 10.52 for mary, 9.52 for cole.